### v2

# LangSplatV2: High-dimensional 3D Language Gaussian Splatting with 450+ FPS

Abstract
In this paper, we introduce LangSplatV2, which achieves high-dimensional feature splatting at 476.2 FPS and 3D open-vocabulary text querying at 384.6 FPS for high-resolution images, providing a 42 $\times$ speedup and a 47 $\times$ boost over LangSplat respectively, along with improved query accuracy. LangSplat employs Gaussian Splatting to embed 2D CLIP language features into 3D, significantly enhancing speed and learning a precise 3D language field with SAM semantics. Such advancements in 3D language fields are crucial for applications that require language interaction within complex scenes. However, LangSplat does not yet achieve real-time inference performance (8.2 FPS), even with advanced A100 GPUs, severely limiting its broader application. In this paper, we first conduct a detailed time analysis of LangSplat, identifying the heavyweight decoder as the primary speed bottleneck. Our solution, LangSplatV2 assumes that each Gaussian acts as a sparse code within a global dictionary, leading to the learning of a 3D sparse coefficient field that entirely eliminates the need for a heavyweight decoder. By leveraging this sparsity, we further propose an efficient sparse coefficient splatting method with CUDA optimization, rendering high-dimensional feature maps at high quality while incurring only the time cost of splatting an ultra-low-dimensional feature. Our experimental results demonstrate that LangSplatV2 not only achieves better or competitive query accuracy but is also significantly faster. Codes and demos are available at our project page: https://langsplat-v2.github.io.
footnotetext: Equal contribution.
1 Introduction
Seamless and intuitive interactions between humans and complex 3D environments kobayashi2022decomposing ; tschernezki2022neural are paramount for a wide range of applications, such as augmented reality craig2013understanding ; azuma1997survey and intelligent robotics peng2023openscene ; huang2023visual ; cascante2022simvqa . To achieve such interactions, systems must understand and respond to natural language queries in real time. This capability hinges on advancements in 3D language fields—a technology at the intersection of vision-language models radford2021learning ; liu2024visual and 3D environmental modeling mildenhall2021nerf ; kerbl20233d .
LangSplat qin2023langsplat , a pioneering model in this domain, leverages 3D Gaussian Splatting to embed 2D CLIP language features into 3D spaces. This method has significantly accelerated the 3D query time, achieving speeds up to 199 $\times$ faster than its predecessors kerr2023lerf , which is critical for applications in scenarios where rapid response is essential. Although many recent improvements have been proposed shorinwa2024fast ; peng2024gags , the inference speed remains a major concern, especially at high resolutions, which hinders their widespread application.
Real-time querying is crucial for applications such as real-time navigation huang2023visual , interactive gaming koh2010integrated , and on-the-fly educational tools godwin2016augmented , where delays can disrupt user experience and functionality. To identify the speed bottleneck of LangSplat, we decompose the entire querying process into three stages: rendering, decoding, and post-processing. Then, we provide a detailed time analysis on the LERF dataset with one A100 GPU. Results in Table 1 show that each query takes 122.1 ms for LangSplat. With some simple engineering modifications, the rendering and post-processing time can be significantly reduced, we term this new version of LangSplat as LangSplat*. The performance of LangSplat* revealed that the decoding stage, which is responsible for transforming low-dimensional latent features into high-dimensional CLIP features with a heavy-weight multi-layer perceptron (MLP), is the primary bottleneck. LangSplat introduces the MLP decoder to significantly reduce the training memory and time cost. Consequently, an additional decoding stage with an MLP is inevitable for test-time querying and reducing the dimensionality of high-dimensional features lowers the accuracy of the queries. However, simply removing the decoder and directly training the high-dimensional CLIP feature for each Gaussian dramatically increases the rendering time. As shown in Figure 1, as the feature splatting dimension increases, the rendering time of LangSplat significantly increases. Specifically, rendering a feature with a dimension of 1536 (assuming we render three semantic levels of 512-dimensional CLIP feature fields in parallel) is 15 times slower than rendering a 3-dimensional field on one A100 GPU.
Figure 1: Feature rendering time comparison with different GPUs. Note that the less advanced GPUs (RTX 3090 and RTX 4090) cannot accommodate the LangSplat model with feature dimensions of 1024 or higher due to running out of memory.
To address the decoding bottleneck in LangSplat, we introduce LangSplatV2, which drastically enhances querying speed without sacrificing querying accuracy. As shown in Figure 1, our LangSplatV2 successfully decouples rendering speed from the dimensionality of rendering features, enabling the rendering of high-dimensional features at the computational cost of splatting an ultra-low-dimensional feature. As the millions of 3D Gaussian points actually represent a limited number of unique semantics for a 3D scene, our LangSplatV2 assumes that each Gaussian can be represented as a sparse coding over a group of global basis vectors. Then, we derive that learning a high-dimensional feature field is equivalent to learning a 3D coefficient field and a global codebook. For each 3D Gaussian point, instead of augmenting a language feature, we learn sparse coefficients. During testing, we utilize the sparsity and propose an efficient sparse coefficient splatting method with CUDA optimization, which effectively decouples the rendering dimension with feature dimensions. As our method removes the autoencoder and directly learns the 3D field from high-dimensional 2D CLIP features, more accurate language fields are also obtained. Experimental results demonstrate that our LangSplatV2 not only achieves higher accuracy but also delivers a substantial speedup: 42 $\times$ faster than LangSplat for high-dimensional feature rendering and 47 $\times$ faster for open-vocabulary 3D querying.
2 Related Work
3D Gaussian Splatting. Recently, 3D Gaussian Splatting (3D-GS) kerbl20233d has received huge attention yu2024mip ; fu2024colmap due to its significant speed advantage over neural radiance fields (NeRFs) mildenhall2021nerf . As a fundamental 3D modeling technology, it’s been widely used in many downstream tasks yu2023cogs ; fan2024momentum ; liu2023humangaussian ; lu20243d ; han2024reparo . Yang et al. yang2023deformable extended 3D Gaussian Splatting to dynamic scenes by modeling deformable 3D Gaussians. Concurrently, 4D Gaussian Splatting wu20234d proposes to use multi-resolution voxel planes to model deformed Gaussians and attains high-quality video rendering results in real-time. Due to the explicit 3D modeling nature of 3D Gaussian Splatting, it has also been used for 3D editing. Gaussianeditor chen2023gaussianeditor offers swift and controllable 3D editing by tracing the editing target throughout the training process. Another concurrent work fang2023gaussianeditor edits 3D scenes using text instruction, which attains two times faster training speed compared with Instruct-NeRF2NeRF haque2023instruct . Many methods zhou2024dreamscene360 ; tang2024lgm ; ren2023dreamgaussian4d utilize 3D Gaussian Splatting for text-driven 3D generation. GSGEN chen2023text incorporates direct geometric priors from 3D Gaussians and obtains a text-to-3D generation model that captures high-frequency components. 3D-GS has also been adopted for generative 3D reconstruction, which has recently attracted significant interest owing to its effectiveness with sparse observations. The first framework of its kind was introduced by Liu et al. liu2024reconx , who pioneered generative 3D reconstruction by modeling it as a temporal generation task. Wang et al. wang2025videoscene subsequently accelerated this approach by using one-step diffusion. Recently, 3D language fields have also embraced 3D-GS. LangSplat qin2023langsplat adopts 3D-GS to model a 3D language field and is 199 $\times$ faster than LERF kerr2023lerf . However, LangSplat is still not a real-time method, which undermines its broad application prospects.
3D Language Field. The concept of 3D language fields shen2023distilled ; siddiqui2023panoptic ; zhi2021place ; li20254d has emerged as an intersection of vision-language models and 3D modeling, aiming to create interactive environments that can be manipulated and queried using natural language. LERF kerr2023lerf explores 3D open-vocabulary text queries by embedding the CLIP feature into a 3D field. Liu et al. liu2023weakly also distilling knowledge from pre-trained foundation models like CLIP radford2021learning and DINO caron2021emerging into NeRF mildenhall2021nerf . It further proposes relevancy-distribution alignment and feature-distribution alignment losses to address the CLIP feature ambiguity issue. LangSplat improves LERF with 3D Gaussian Splatting and attains a significant speedup. There are also some other works shi2024language ; zuo2024fmgs ; zhou2023feature ; li2024langsurf employing 3D Gaussian Splatting for 3D language field modeling, they usually suffer from imprecise language fields like LERF, which has been well addressed by LangSplat. Following LangSplat, GOI qu2024goi3dgaussiansoptimizable proposes simplifying feature selection by dividing the feature space with a hyperplane, keeping only the features most relevant to the query. GAGS peng2024gags enhances multiview consistency by linking SAM’s prompt point density with camera distances and introducing an unsupervised granularity factor to selectively distill consistent 2D features. However, these methods both use a decoder to map the low-dimensional feature to high-dimensional space, which is a primary bottleneck of real-time query as shown in section 3.2.
3D Gaussian Splatting Acceleration. To reduce computational overhead and improve rendering speed, recent works have proposed various strategies to accelerate 3D-GS. One line of work compresses the Gaussian representation using vector quantization and codebooks. For instance, Compressed 3D Gaussian Splatting niedermayr2024compressed adopts sensitivity-aware clustering and quantization-aware training to compress both directional colors and Gaussian parameters, achieving up to $31\times$ compression and $4\times$ rendering speedup. CompGS navaneet2024compgs similarly applies K-means-based quantization to achieve over $40\times$ storage reduction with minimal quality loss. Other approaches focus on structural regularization or architectural simplification. Self-Organizing Gaussian Grids morgenstern2024compact reorganize Gaussian parameters into spatially coherent 2D grids, reducing redundancy via local homogeneity. LightGaussian fan2024lightgaussian uses pruning based on global significance and vector quantization, achieving $15\times$ compression and 200+ FPS rendering. C3DGS lee2024compact proposes to reduce the number of Gaussians via a learnable mask and introduces compact view-dependent color encoding, achieving a $25\times$ reduction in storage. Techniques like Speedy-Splat hanson2025speedy and SORT-Free Splatting hou2024sort improve runtime performance through optimized rasterization and differentiable approximations of alpha blending. In contrast to these works that focus on accelerating RGB-based scene rendering, our method tackles the unique challenge of handling high-dimensional Gaussian features. Extending prior compression techniques niedermayr2024compressed ; fan2024lightgaussian to our setting would require learning and quantizing such high-dimensional attributes per Gaussian, leading to excessive memory usage. Moreover, directly rendering with full 512D features remains computationally demanding. In contrast, our method circumvents the need to directly learn and render high-dimensional features for 3D Gaussian points.
3 Proposed Approach
3.1 Preliminaries
3D Gaussian Splatting employs point-based rendering technologies pfister2000surfels ; zwicker2001surface and models the scene geometry as a set of 3D Gaussian points. Each Gaussian point is associated with the following attributes: Gaussian center position $\mu\in\mathbb{R}^{3}$, a covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$, color described $c\in\mathbb{R}^{3}$ by spherical harmonic (SH) coefficients, and opacity $o\in\mathbb{R}$. Each Gaussian point is represented as:
$G(x)=\exp(-\frac{1}{2}(x-\mu)^{\top}\Sigma^{-1}(x-\mu)).$ (1)
The covariance matrix $\Sigma$ is further represented with a rotation matrix and a scaling factor matrix:
$\Sigma=RSS^{\top}R^{\top},$ (2)
where $R\in\mathbb{R}^{4}$ denotes the rotation matrix and $S\in\mathbb{R}^{3}$ represents the scaling matrix.
Based on EWA volume splatting zwicker2001ewa , 3D Gaussian Splatting projects the 3D Gaussian points onto a 2D image plane, which blends the ordered Gaussian points that overlap with the rendered pixel $v$:
$C(v)=\sum_{i\in\mathcal{N}}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (3)
where $\mathcal{N}$ represents the ordered Gaussian points within a tile, and $\alpha_{i}=o_{i}G^{2D}_{i}(v)$. The $G^{2D}_{i}(\cdot)$ means the projected 2D Gaussian distribution of the $i$-th 3D Gaussian point.
As 3D Gaussian Splatting exhibits excellent real-time rendering speed while maintaining high rendering quality even at 1080p resolution, many methods wu20234d ; yang2023deformable ; zou2023triplane have adopted it as a 3D scene modeling technology to accelerate the rendering speed. LangSplat aims to build a 3D language field to support open-vocabulary querying within 3D spaces. It extends each 3D Gaussian with a language embedding and supervises the 3D language Gaussians with 2D CLIP image features. LangSplat presents two main contributions to make it faster and more accurate. First, instead of using multiple patch-wise CLIP features with varying scales, which leads to imprecise and vague 3D language fields, LangSplat employs the CLIP features of SAM masks with three pre-defined SAM hierarchical semantic scales to obtain precise language fields with clear boundaries. Second, as CLIP embeddings are high-dimensional, directly splatting at CLIP feature space poses huge challenges in training memory and time cost. Therefore, LangSplat proposes a scene-specific autoencoder, which compresses the high-dimensional (512-D) CLIP features into low-dimensional (3-D) latent features. The rendering process of LangSplat follows:
$\bm{F}(v)=\sum_{i\in\mathcal{N}}\bm{f}_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (4)
where $\bm{f}_{i}\in\mathbb{R}^{d}$ represents the augmented $d$-dimensional latent feature at $i$-th 3D Gaussian point and $\bm{F}(v)$ denotes the rendered latent feature at pixel $v$. After obtaining the rendered feature, LangSplat uses the decoder $g_{d}(\bm{F}(v))\in\mathbb{R}^{D}$ from the trained autoencoder $g$ to decode the latent feature back to CLIP feature space. In the end, the LangSplat attains an accurate 3D language field while being 199 $\times$ faster than the previous state-of-the-art method LERF kerr2023lerf .
3.2 Bottleneck Analysis
While LangSplat achieved significant speedup over other methods, it still cannot perform real-time open-vocabulary 3D querying at high resolution. However, there is a high demand for real-time 3D querying for many applications, such as Augmented Reality (AR) jain2023real , intelligent robotics kerr2023lerf .
Table 1: The stage-wise time cost (ms) for LangSplat and our improvements on the LERF dataset with one A100 GPU. LangSplat* is modified from LangSplat with simple engineering optimization. LangSplatV2 achieves a speed of up to 384.6 FPS for open-vocabulary 3D querying.
Method Rendering Decoding Post-Processing Total Speed (FPS)
LangSplat 6.0 83.1 33.0 122.1 8.2
LangSplat* 2.0 83.1 0.5 85.6 11.7
LangSplatV2 2.0 0.1 0.5 2.6 384.6
To further improve the query speed of LangSplat, we first analyze each step of the query process and identify the bottleneck. For a given text query, the inference of LangSplat can be divided into three stages: rendering, decoding, and post-processing. The rendering stage will render the learned 3D language Gaussians into a 2D image plane and get a rendered latent feature map $\bm{F}\in\mathbb{R}^{H\times W\times d}$, where $H,W$ denote the height and width of the 2D image size, respectively. As the rendered feature map only encodes language features in the low-dimensional latent space, a decoding stage is followed to obtain the feature map at CLIP space with a decoder $g_{d}(\bm{F})\in\mathbb{R}^{H\times W\times D}$, where the decoder $g_{d}(\cdot)$ is implemented with an MLP. The last post-processing stage computes the relevancy score with the obtained $D$-dimensional feature following LERF kerr2023lerf and remove some noise in the relevancy score map. Specifically, a mean filter is applied to the relevancy score map. Note that LangSplat adopts the three semantic levels introduced by SAM kirillov2023segment , so the above operations are repeated three times for three SAM semantic levels and the post-processing stage needs to select one semantic level for prediction with some specific strategies qin2023langsplat .
We test the stage-wise inference time of LangSplat on the LERF dataset with one A100 GPU. The results in Table 1 show that the rendering stage takes 6.0 ms, the decoding stage takes 83.1 ms, and the post-processing stage takes 33.0 ms. While all stages occupy a considerable amount of time, we can significantly improve them through two simple modifications. First, we notice that LangSplat implements the post-processing stage on the CPU, which could be slow when it involves mean filter operations. We implement the post-processing stage on the GPU and observe that the time cost of the post-processing stage is now negligible compared with other stages. The second modification is scale parallelization. LangSplat performs text querying on three semantic levels sequentially, meaning the rendering stage is performed three times separately. However, we can perform inference at three semantic scales in parallel. For example, instead of splatting $d$-dimensional features three times, we directly splat $3d$-dimensional features for once. We use LangSplat* to denote the version with these simple modifications. We observe that for LangSplat*, the rendering stage takes 2.0 ms, and the post-processing stage takes 0.5 ms. In these three stages, the decoding phase occupies 97.1% of the total query time, making it the primary bottleneck in achieving real-time 3D querying.
Figure 2: The framework of LangSplatV2. LangSplatV2 introduces a sparse coefficient for each Gaussian point and a shared global codebook for the entire scene.
3.3 3D Sparse Coefficient Field
Different from 3D Gaussian Splatting, which renders 3-dimensional RGB color, 3D language fields need to splat high-dimensional features. Directly splatting high-dimensional features significantly decreases the rendering speed, as shown in Figure 1. To address this issue, existing methods usually model $d$-dimensional ($d<<D$) latent features in 3D followed by either an online decoder zhou2023feature or an offline decoder qin2023langsplat to decode latent features back to $D$-dimensional features in 2D image space. The high dimension gap between latent features and decoded features implies a heavyweight MLP is required to ensure the accuracy of the decoding stage, which becomes the primary speed bottleneck.
In LangSplatV2, instead of splatting the $d$-dimensional latent language features, we propose to model a 3D sparse coefficient field. As shown in Eq. 4, LangSplat assigns each Gaussian point with a language feature $\bm{f}_{i}$, which represents the semantics associated with the Gaussian point. As LangSplat could create millions of Gaussian points, there will be millions of unique language features. However, this is highly inefficient as the unique semantics within a scene are quite limited and much smaller than the number of Gaussian points. As a matter of fact, many Gaussian points share similar semantics. Therefore, we assume the language embedding of every Gaussian point within a scene can be represented as a sparse coding of $L$ global basis vectors $\mathcal{S}=[\bm{s}_{1},\bm{s}_{2},...,\bm{s}_{L}]^{\top}\in\mathbb{R}^{L\times D}$. These $L$ basis vectors serve as the global codebook and each Gaussian point is computed by linearly combining a small number of local basis vectors. We assume that only $K$ ($K<<L$) basis vectors from $L$ global codebook are used to represent the language embedding of a Gaussian point. Then we define $\bm{w}_{i}=[w_{i,1},w_{i,2},...,w_{i,L}]^{\top}\in\mathbb{R}^{1\times L}$ as the associated sparse coefficients for $i$-th Gaussian point, where $\sum_{l=1}^{L}w_{i,l}=1$ and only $K$ elements are non-zero values while all other elements are zeros. With slightly abuse of the notion $\bm{f}_{i}$, the language embedding $\bm{f}_{i}\in\mathbb{R}^{D}$ of $i$-th Gaussian point is represented as:
$\bm{f}_{i}=\bm{w}_{i}\mathcal{S}=\sum_{l=1}^{L}w_{i,l}\bm{s}_{l}.$ (5)
If we directly render the $D$-dimensional language field without compressing CLIP features, we have:
$\bm{S}=\sum_{i\in\mathcal{N}}\bm{w}_{i}\mathcal{S}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (6)
where $\bm{S}$ is the rendered $D$-dimensional CLIP feature. We set $e_{i}=\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$, then we have:
$\bm{S}=\sum_{i\in\mathcal{N}}\bm{w}_{i}\mathcal{S}e_{i}=(\sum_{i\in\mathcal{N}}e_{i}\bm{w}_{i})\mathcal{S}.$ (7)
Eq. 7 shows that rendering the $D$-dimensional CLIP features is equivalent to first rendering the sparse coefficients $\bm{w}(i)$, and then performing a matrix multiplication with the global dictionary $\mathcal{S}$.
Therefore, we propose to learn a $L$-dimensional sparse coefficient for each Gaussian point and a global codebook with a size of $L$ basis vectors. To learn a sparse distribution for the coefficient $\bm{w}_{i}$, we first apply a softmax function to normalize the $L$-dimensional parameter, then preserve the top-$K$ values as non-zeros elements and set the remaining elements to zeros. We will re-normalize the top-$K$ values with their sum to ensure the sum of the coefficient equals one. The $L$-dimensional 3D sparse coefficient field and the $L$ basis vectors are jointly learned. Figure 2 visualizes the framework.
Compared to LangSplat, LangSplatV2 requires only a simple matrix multiplication after rendering the weight map, instead of relying on a heavyweight decoder, thus overcoming LangSplat’s main inference speed bottleneck. Furthermore, the $D$-dimensional global codebook eliminates reconstruction loss from dimensionality reduction, enabling better modeling of high-dimensional features in the CLIP space and improving query accuracy.
3.4 Efficient Sparse Coefficient Splatting
By learning a 3D sparse coefficient field, the MLP decoder is entirely removed, eliminating the associated computational overhead of the MLP. However, we still need to perform a $3L$-dimensional feature rendering and a matrix multiplication. Our experiments show that the time for matrix multiplication (in Table 1, listed as the decoding stage of LangSplatV2) is negligible compared to the rendering process. However, rendering high-dimensional features with dimensionality $L$ remains computationally demanding. To address this issue, we propose an efficient sparse coefficient splatting method with CUDA optimization. By exploiting the sparsity of the coefficient field, we can achieve $L$-dimensional feature rendering at the cost of only $K$-dimensional rendering, where $K\ll L$.
Figure 3: Our efficient sparse coefficient splatting method accelerates the speed of alpha-blending by utilizing the property of the learned sparse coefficient field and neglecting zero elements.
In the CUDA implementation of 3D Gaussian Splatting and LangSplat, each thread performs alpha-blending of the $|\mathcal{N}|$ ordered Gaussian points within a tile. With an $L$-dimensional rendering, each thread sequentially computes $L$ channels, leading to a computational complexity of $O(|\mathcal{N}|L)$. When $L$ becomes sufficiently large, this alpha-blending computation becomes the key bottleneck and significantly increases the overall computational overhead as $L$ grows, as shown in Figure 1.
To mitigate this, we utilize the sparse nature of the learned coefficient field during test-time querying, as shown in Figure 3. Although each Gaussian point has an $L$-dimensional coefficient, only $K$ dimensions contain non-zero values. Therefore we can only perform alpha-blending for the non-zero elements, as the blending of zero elements does not alter the result. This effectively reduces the computational complexity from $O(|\mathcal{N}|L)$ to $O(|\mathcal{N}|K)$, with $K$ significantly smaller than $L$. In practice, we set $K=4$, which allows us to simultaneously render three semantic scales with an effective feature rendering dimension in CUDA of only $12$, yielding a high-quality feature map of $1,536$ dimensions. It makes the rendering highly efficient without compromising feature quality.
Specifically, each Gaussian’s $L$-dimensional sparse coefficient $\bm{w}_{i}$ can be fully represented by its top-$K$ non-zero elements. We store these as two $K$-dimensional arrays for each Gaussian: top-$K$ indices and top-$K$ coefficients. The top-$K$ indices are the positions of the top-$K$ non-zero elements within the $L$-dimensional vector and the top-$K$ coefficients are the values of these top-$K$ non-zero elements. During rendering, each CUDA thread performs alpha-blending only for the $K$ non-zero coefficients. For each Gaussian point within the tile, the CUDA thread will access the indices and coefficients of the top-$K$ elements and perform weighted summation solely on these $K$-dimensional indices, avoiding computation on zero elements. In this way, LangSplatV2 can obtain high-dimensional ($D$) feature splatting results at the cost of only ultra-low-dimensional ($K$) feature splatting.
4 Experiments
4.1 Datasets and Details
Datasets. We evaluate our method on the LERF, 3D-OVS, and Mip-NeRF360 datasets. The LERF dataset kerr2023lerf , captured using the iPhone App Polycam, contains in-the-wild scenes. For the open-vocabulary 3D object localization task, we adopt the augmented localization annotations provided by LangSplat qin2023langsplat on the LERF dataset. Additionally, we use the segmentation ground truth from LangSplat qin2023langsplat for the open-vocabulary 3D segmentation task on LERF. Beyond LERF, we also conduct 3D segmentation experiments on the 3D-OVS and Mip-NeRF360 JTBarron2022mip datasets. The 3D-OVS dataset liu2023weakly includes a collection of long-tail objects captured in diverse poses and backgrounds. Moreover, we evaluate our method on Mip-NeRF360, which consists of multi-view indoor and outdoor scene images, with segmentation labels annotated by GAGS peng2024gags . For evaluation, we use localization accuracy for the 3D object localization task and report the average IoU scores for the open-vocabulary 3D segmentation task.
Implementation Details. Following LangSplat qin2023langsplat , we use the OpenCLIP ViT-B/16 model to extract CLIP features. We employ the ViT-H model for SAM kirillov2023segment to segment images and obtain masks with three hierarchical semantics. The codebook size $L$ is set to 64 and the $K$ is set to 4. During test-time querying, we render three semantic scales simultaneously, leading to the actual rendering dimension of 12. The 3D Gaussians are first trained with RGB supervision for 30,000 iterations to reconstruct the RGB scene. Then we train another 10,000 iterations for the 3D sparse coefficient field by fixing all other 3D Gaussian parameters. All our experiments are conducted on one A100 GPU.
Table 2: Quantitative comparisons of open-vocabulary 3D object localization and 3D semantic segmentation on the LERF dataset. We report the mean accuracy (%) and the mean IoU scores (%).
Method 3D Object Localization 3D Semantic Segmentation
Ramen Teatime Kitchen Figurines Overall Ramen Teatime Kitchen Figurines Overall
GS-Grouping ye2024gsgroup 32.4 69.5 50.0 44.6 49.1 26.4 54.0 31.3 34.6 36.6
LEGaussian shi2024language 69.0 79.7 63.6 57.1 67.4 20.2 32.3 22.3 23.4 24.6
GOI qu2024goi3dgaussiansoptimizable 56.3 67.8 68.2 44.6 59.2 33.7 55.8 54.5 23.9 42.0
GAGS peng2024gags 69.0 88.1 90.9 78.6 81.7 46.8 60.3 55.8 53.6 54.1
LangSplat qin2023langsplat 73.2 88.1 95.5 80.4 84.3 51.2 65.1 44.5 44.7 51.4
LangSplatV2 74.7 93.2 86.4 82.1 84.1 51.8 72.2 59.1 56.4 59.9
4.2 Quantitative Results
Figure 4: Qualitative comparisons of open-vocabulary 3D object localization on the LERF dataset. The red points are the model predictions and the black dashed bounding boxes denote the annotations. We observe that LangSplatV2 generates better results than LangSplat.
Time Analysis. To assess the speed improvement of LangSplatV2 over its LangSplat, we conducted a detailed time analysis using the LERF dataset, with the computations performed on a single A100 GPU. Table 1 presents a stage-wise breakdown of the time costs for LangSplat, LangSplat*, and LangSplatV2. LangSplat, our initial model, achieves a frame rate of 8.2 FPS. The LangSplat* improvements mainly focused on rendering and post-processing optimizations, which reduced the total time to 85.6 ms and increased the speed to 11.7 FPS. Here, the rendering time was significantly cut to 2.0 ms, and the post-processing time was reduced to less than 1.0 ms with our proposed simple engineering modifications. In contrast, LangSplatV2 proposed an efficient sparse coefficient splatting method, that entirely removed the MLP decoder in the decoding stage. The only operation in the decoding stage is performing a matrix multiplication between the rendered $L$-dimensional coefficient and the global dictionary, which takes only 0.1 ms. While completely removing the MLP decoder, our LangSpaltV2 only needs 2.0 ms to render the sparse coefficient, leading to an overall 476.2 FPS for obtaining high-dimensional feature maps. In the end, LangSplatV2 achieved an exceptional reduction in total querying time to 2.6 ms. These results not only demonstrate a substantial enhancement in processing speed but also affirm the real-time capabilities of LangSplatV2 for high-resolution, complex 3D scene querying.
Table 3: Quantitative 3D semantic segmentation results on 3D-OVS, reported as mean IoU (%).
Method Bed Bench Room Sofa Lawn Overall
FFD kobayashi2022decomposing 56.6 6.1 25.1 3.7 42.9 26.9
LERF kerr2023lerf 73.5 53.2 46.6 27.0 73.7 54.8
3D-OVS liu2023weakly 89.5 89.3 92.8 74.0 88.2 86.8
GS-Grouping ye2024gsgroup 83.0 91.5 85.9 87.3 90.6 87.7
LEGaussian shi2024language 84.9 91.1 86.0 87.8 92.5 88.5
GOI qu2024goi3dgaussiansoptimizable 89.4 92.8 91.3 85.6 94.1 90.6
LangSplat qin2023langsplat 92.5 94.2 94.1 90.0 96.1 93.4
LangSplatV2 93.0 94.9 96.1 92.3 96.6 94.6
Main Results on the LERF dataset. Table 2 demonstrates the comparisons on the LERF dataset. For the open-vocabulary 3D object localization task, LangSplatV2 achieved the highest accuracy with notable improvements over previous methods for most scenes, including “ramen", “figurines", and “teatime". In the "Kitchen" scene, LangSplat achieved a higher accuracy than LangSplatV2. This difference is mainly due to the limited sample size (22 examples in total), with only a two-prediction gap (21 vs. 19). Given LangSplat’s high variance in this scene, we ran it multiple times, obtaining an average of 17.5 $\pm$ 1.8 correct predictions, which suggests that LangSplatV2 provides more consistent performance advantage across different scenarios. Results on 3D semantic segmentation demonstrate that LangSplatV2 outperforms all other methods, and it consistently surpasses LangSplat across all scenes, highlighting the effectiveness of our proposed method.
Table 4: Quantitative 3D semantic segmentation results on Mip-NeRF360, reported as mean IoU (%).
Method Room Counter Garden Bonsai Overall
GS-Grouping ye2024gsgroup 54.4 47.7 40.4 54.1 49.2
LEGaussian shi2024language 25.5 35.3 33.2 22.3 29.1
GOI qu2024goi3dgaussiansoptimizable 60.3 46.6 59.8 67.3 58.5
GAGS peng2024gags 65.2 61.1 61.2 70.5 64.5
LangSplat qin2023langsplat 53.2 68.8 51.9 55.4 57.3
LangSplatV2 64.3 75.1 65.0 73.1 69.4
Main Results on the 3D-OVS dataset. Table 3 shows the quantitative comparisons of open-vocabulary 3D semantic segmentation performance, on the 3D-OVS dataset across five test scenes. We see that LangSplatV2 achieves the highest overall mean IoU score of 94.6%. It significantly outperforms LangSplat across all different scenes, further validating its ability to build more accurate language fields.
Main Results on the Mip-NeRF360 dataset. Table 4 presents the comparison results on the Mip-NeRF360 dataset. Notably, LangSplatV2 achieved a significant improvement in the overall average IoU score, reaching 69.4%, which is a marked advancement over LangSplat’s 57.3%.
Table 5: Ablation study on codebook size $L$. We report the average performance of the 3D object localization and 3D semantic segmentation tasks on the LERF dataset.
$L$ 32 64 128
Accuracy (%) 72.8 84.1 84.1
IoU (%) 53.9 59.9 60.5
Table 6: Ablation study on different $K$. We report the average performance of the 3D object localization and 3D semantic segmentation tasks on the LERF dataset.
$K$ 2 4 6 8
Accuracy (%) 79.4 84.1 84.4 83.6
IoU (%) 54.4 59.9 59.9 59.9
Ablation Study. In Table 6, we evaluate the effect of codebook sizes $L$ on the localization and segmentation tasks. We report the mean accuracy and IoU scores on the LERF dataset. Our results reveal that increasing L from 32 to 64 yields a notable improvement in both localization accuracy (72.8% to 84.1%) and segmentation IoU (53.9% to 59.9%). This suggests that a larger codebook size better captures the semantic fields in complex scenes, allowing for more precise language representations. However, further increasing $L$ results in only slightly increased performance, suggesting that $L=64$ has already reached saturation on the LERF dataset. Table 6 further reports the effects of different $K$. We observe that $K=4$ can saturate the performance of the Gaussian model, and a larger K will increase the rendering dimensions, thereby slowing down the rendering speed. More ablation studies and details can be found in the supplementary material.
Figure 5: Qualitative comparisons of open-vocabulary 3D semantic segmentation on the LERF, Mip-NeRF360 and 3D-OVS dataset. We can see that our LangSplatV2 generates better masks than LangSplat, which shows the effectiveness of our LangSplatV2.
Table 7: Quantitative comparisons of training cost on the LERF dataset.
Method LangSplat qin2023langsplat LEGaussian shi2024language LangSplatV2 LangSplat-512D qin2023langsplat
Training Time (h) 1.0 1.3 3.0 45.0
Training Memory (GB) 6.2 11 21.2 47.4
[…]
Table 8: Quantitative comparisons with LEGaussian on three benchmarks.
Method Segmentation IoU (%) Query Time (ms) GPU Memory (GB)
LERF 3DOVS Mip. LERF 3DOVS Mip. LERF 3DOVS Mip.
LEGaussian shi2024language 24.6 88.5 29.1 36.7 59.6 58.0 8.2 10.5 14.0
LangSplatV2 59.9 94.6 69.4 2.6 4.8 3.8 7.2 11.6 16.9
Discussion. Recent work has also explored codebook-based language field representations, such as LEGaussians shi2024language . However, our approach departs significantly from theirs in several key aspects. First, LEGaussians construct a 2D feature codebook from 2D images while LangSplatV2 directly learns a global codebook in 3D space, which is more efficient. Second, LEGaussians still rely on an MLP to compress features and do not exploit sparsity. In contrast, our method completely removes the MLP and further enforces and leverages sparsity in both representation and rendering. As shown in Table 8, our method achieves superior semantic segmentation accuracy, faster inference speed, and comparable memory usage across three benchmark datasets, which demonstrates the effectiveness of our approach. For a more detailed comparison and discussion, please refer to the Appendix.
4.3 Qualitative Results
Figure 4 visualizes the open-vocabulary 3D object localization results. We observe that our LangSpaltV2 can give more accurate localization predictions compared with LangSplat. For example, our LangSplatV2 can give the correct prediction for “pumpkin" while LangSplat entirely fails in this query. Figure 5 visualizes the open-vocabulary 3D segmentation results in three datasets. We observe that our LangSplatV2 can generate more accurate masks compared with LangSplat. For example, in the “Kitchen" scene, LangSplat predicts a noisy mask for the “ketchup" query, while our LangSplatV2 generates more precise and clean masks.
5 Conclusion
Conclusion. In this paper, we have presented LangSplatV2 for high-dimensional language feature splatting in 3D open-vocabulary querying, designed to overcome the speed limitations of LangSplat. Our analysis identified the decoding stage as the primary bottleneck in LangSplat. By modeling each Gaussian point as a sparse code over a global dictionary, LangSplatV2 removes the need for a heavyweight decoder. We further proposed an efficient sparse coefficient splatting method with CUDA optimization, allowing LangSplatV2 to achieve high-dimensional feature splatting results at the cost of splatting ultra-low-dimensional features.
Limitations and broader impacts. While LangSplatV2 demonstrates improved performance and faster inference compared to LangSplat, it comes with increased training cost due to the need to construct high-dimensional semantic fields during training. Furthermore, as our method directly inherits the semantic representations from the CLIP model, it also inherits its inherent biases. Addressing such biases remains an open research challenge and may benefit from recent advances in fairness-aware versions of CLIP.
Acknowledgment
This research is supported in part by the NIH grant R01HD104969.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discussed the limitations in the Conclusion Section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This paper does not involve theoretical contributions.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We have included an Algorithm to demonstrate how to reproduce our method. We have released the code.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We have released the source codes.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have included it in the Experiments Section.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A]
  34. Justification: It is not included in all previous work in this field.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We have included it in the Experiments Section.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our research conducted in the paper conforms to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We have discussed the broader impact in the Conclusion Section.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have properly cited the original paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Algorithm
The proposed efficient sparse coefficient splatting process is shown in Algorithm 1.
Algorithm 1 Efficient Sparse Coefficient Splatting
 Initialize $\bm{W}\in\mathbb{R}^{L}$ to zero
for Gaussian point $i$ $\in\mathcal{N}$ do
  Retrieve top-$K$ indices $\{j_{1},j_{2},...,j_{K}\}$ and corresponding coefficients $\{w_{i,j_{1}},w_{i,j_{2}},...,w_{i,j_{K}}\}$
  for each $j\in\{j_{1},...,j_{K}\}$ do
   $\bm{W}[j]\mathrel{+}=w_{i,j}\cdot\alpha_{i}\cdot\prod_{m=1}^{i-1}(1-\alpha_{m})$
  end for
end for
Appendix B Discussion
LangSplatV2 proposes to model the high-dimensional language features of each Gaussian point as a sparse code within a global dictionary. While similar concepts, such as quantizing Gaussian point attributes (e.g., SH coefficients fan2024lightgaussian , opacity niedermayr2024compressed , and semantic features shi2024language ) into a codebook, have been explored in the domains of compression niedermayr2024compressed and language embedding shi2024language , our method is fundamentally different.
In the compression domain, existing approaches niedermayr2024compressed ; fan2024lightgaussian first learn unique attributes for each Gaussian point, which are then quantized into a codebook using a quantizer. Extending this concept to our scenario would require learning 512-dimensional semantic features for each Gaussian point before quantization. This process significantly increases memory consumption and training time, often exceeding the capacity of GPUs like the NVIDIA 3090 or 4090, leading to out-of-memory (OOM) errors. Moreover, even with advanced GPUs that can handle the training, the rendering process would still involve managing 512-dimensional features. As shown in Figure 1, this dramatically reduces rendering speed compared to LangSplatV2.
In the context of language embedding, LEGaussians shi2024language adopt a codebook-based approach to accelerate training. However, our method is significantly different from LEGaussians in three key aspects: 1) Global 3D Codebook vs. 2D Codebook: LEGaussians first train a 2D codebook by quantizing features in 2D images, then learn a 3D model to predict the one-hot class category indicating the index of the codebook. In contrast, LangSplatV2 learns a global 3D codebook shared among all 3D Gaussian points. By avoiding the additional reconstruction errors caused by first quantizing features in the 2D image plane, our approach reduces overall reconstruction error and more efficiently utilizes the codebook in 3D space. 2) Eliminating the MLP Bottleneck: LEGaussians rely on an MLP to project each Gaussian point’s features into a one-hot class category representing the index of the codebook. Our method removes the need for an MLP entirely, thereby eliminating the speed bottleneck and improving computational efficiency. 3) Efficient Sparse Coefficient Splatting: LangSplatV2 introduces Efficient Sparse Coefficient Splatting, which reduces the effective rendering dimensions to $K$. In contrast, LEGaussians renders with the full feature dimensions of Gaussian points, which inherently slows down the rendering process.
Figure 6: More qualitative comparisons of open-vocabulary 3D object localization on the LERF and Mip-NeRF360 datasets. The red points are the model predictions and the black dashed bounding boxes denote the annotations. We observe that LangSplatV2 generates better results than LangSplat.
Figure 7: More qualitative comparisons of open-vocabulary 3D semantic segmentation on the LERF, Mip-NeRF360 and 3D-OVS dataset. We can see that our LangSplatV2 generates better masks than LangSplat, which shows the effectiveness of our LangSplatV2.
Furthermore, due to the reliance on an MLP decoder, LEGaussians still operate in a lower-dimensional feature space, which limits the expressiveness of its learned representations. Additionally, its two-stage 2D codebook approach fails to effectively incorporate 3D priors, resulting in a suboptimal learned field. In contrast, LangSplatV2 directly models high-dimensional features within a global 3D codebook, effectively capturing the underlying 3D structures. As evidenced by the results in Table 8, our approach outperforms LEGaussians across multiple benchmarks while being significantly faster, demonstrating superior scene reconstruction quality and more accurate language-grounded 3D representations.
To conclude, LangSplatV2 demonstrates a significant leap forward in the efficiency and scalability of modeling high-dimensional language features in 3D scenes. By leveraging the proposed 3D sparse coefficient fields and efficient sparse coefficient splatting techniques, our method reduces computational overhead while maintaining high fidelity in 3D language scene rendering.
Appendix C Evaluation Details
After calculating the similarity between the rendered semantic features and the CLIP features of the query text, we obtained relevance maps $M$ at different semantic levels. Then, we use different strategies for different datasets to obtain the corresponding query results.
LERF. For object localization, we first apply average pooling to smooth $M$ and get $Avg(M)$. Then we select the point with the highest relevancy in $Avg(M)$ as the predicted point. Additionally, we choose the semantic level corresponding to the highest relevancy value among the three levels as the selected level. For semantic segmentation, after getting $Avg(M)$, we select the level in the same way as object localization. We then normalize $Avg(M)$ and use a threshold to obtain a binary mask as the final prediction mask.
3D-OVS. The post-processing pipeline for semantic segmentation is similar to that used for the LERF dataset. Differently, we obtain the smoothed relevancy map $Avg^{\prime}(M)$ as $Avg^{\prime}(M)=0.5\times(M+Avg(M))$, and the level selection is based on the average relevancy within the predicted mask. Additionally, when evaluating query accuracy, we use two coarse-grained semantic levels (whole and part) owing to the simplicity of the scenes in this dataset.
Mip-NeRF360. Similar to the post-processing for the 3D-OVS dataset, we obtain the smoothed relevancy map $Avg^{\prime}(M)$ and select the semantic level based on the average relevancy score. When evaluating query accuracy, we use all three semantic levels, as the complex scenes in this dataset demand finer-grained language feature rendering.
Appendix D More Ablation Study
Time Analysis. We conduct ablation studies on the LERF kerr2023lerf dataset and report the render and decode speed (ms per query) in Table 9. We test the speed on one A100 GPU. The baseline is LangSplat, which renders three-level 3-dimensional language features separately. Rendering three semantic levels in parallel can reduce the rendering time from 6.0 ms to 2.0 ms per query. The sparse coefficient field can significantly speed up the decoding stage from 83.1 ms/q to 0.1 ms/q by changing the heavy weight decoder to a simple matrix multiplication operation. Efficient sparse coefficient splatting method effectively transformed the 192-dimensional rendering into 12-dimensional rendering, reducing the rendering time from 5.3 ms to 2.0 ms per query through CUDA optimization.
Table 9: Speed ablation study on the LERF dataset. Parallel means rendering three semantic levels in parallel, Sparse means 3D sparse coefficient field, Efficient means efficient sparse coefficient splatting with CUDA optimization.
Component Speed (ms/q)
Parallel Sparse Efficient render decode total
6.0 83.1 89.1
2.0 83.1 85.1
5.3 0.1 5.4
2.0 0.1 2.1
$L$ and $K$. In Table 10 and Table 11, we show the ablation study results on each scene in the LERF kerr2023lerf dataset. As we can see, for some complex scenarios, such as Kitchen and Figurines, larger $L$ and $K$ can lead to better performance. However, for all scenarios, the model’s performance is already good enough with the settings of $L=64$ and $K=4$, and increasing $L$ and $K$ will lead to increased computational resources and time cost for training and inference. Therefore, we set $L=64$ and $K=4$ to balance performance and efficiency.
Table 10: Ablation study on codebbok size $L$. We report the localization and the segmentation performance on the LERF dataset.
$L$ 32 64 128
Ramen Accuracy (%) 70.4 74.7 74.7
IoU (%) 50.5 51.8 51.4
Teatime Accuracy (%) 84.8 93.2 91.5
IoU (%) 65.7 72.2 69.8
Kitchen Accuracy (%) 68.2 86.4 86.4
IoU (%) 54.2 59.1 63.2
Figurines Accuracy (%) 67.9 82.1 83.9
IoU (%) 45.3 56.4 57.6
Overall Accuracy (%) 72.8 84.1 84.1
IoU (%) 53.9 59.9 60.5
Table 11: Ablation study on different $K$. We report the localization and the segmentation performance on the LERF dataset.
$K$ 2 4 6 8
Ramen Accuracy (%) 71.8 74.7 71.8 73.2
IoU (%) 50.2 51.8 51.4 51.1
Teatime Accuracy (%) 91.5 93.2 91.5 91.5
IoU (%) 68.8 72.2 70.7 70.6
Kitchen Accuracy (%) 77.3 86.4 95.5 90.9
IoU (%) 48.9 59.1 59.8 60.1
Figurines Accuracy (%) 76.8 82.1 78.6 78.6
IoU (%) 49.6 56.4 57.7 57.6
Overall Accuracy (%) 79.4 84.1 84.4 83.6
IoU (%) 54.4 59.9 59.9 59.9
Appendix E More Visualization Results
Open-vocabulary 3D Object Localization. We visualize more examples on the LERF kerr2023lerf and Mip-NeRF360 kerr2023lerf datasets for open-vocabulary 3D object localization in Figure 6.
Open-vocabulary 3D Semantic Segmentation. We visualize more examples on the LERF kerr2023lerf , Mip-NeRF360 JTBarron2022mip and 3D-OVS liu2023weakly datasets for open-vocabulary 3D semantic segmentation in Figure 7.